OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon if each exterior angle measures 72?

OpenStudy (anonymous):

A. 720 B. 1080 C. 360 D. 180 E. 540 F. 900

Directrix (directrix):

The sum of the exterior angles of any regular polygon is 360. 360/72 = 5 which means that there are 5 interior angles of the regular polygon, and thus 5 sides. The polygon is a pentagon. If each exterior angle of the regular polygon has measure 72, then the adjacent interior angle has measure 180 - 72 = 108. 108 *5 = ? degrees for the sum of the interior anlges of the pentagon. @Danny_Boy Note: Regular polyons are both equiangular and equilateral.

OpenStudy (anonymous):

so my answer would be 180??? @Directrix

Directrix (directrix):

@Danny_Boy >>so my answer would be 180??? NO Multiply 108 times 5 and that is your answer.